### v1

# Seewo’s Submission to MLC-SLM: Lessons learned from Speech Reasoning Language Models

Abstract
This paper presents Seewo’s systems for both tracks of the Multilingual Conversational Speech Language Model Challenge (MLC-SLM), addressing automatic speech recognition (ASR) and speaker diarization with ASR (SD-ASR). We introduce a multi-stage training pipeline that explicitly enhances reasoning and self-correction in speech language models for ASR. Our approach combines curriculum learning for progressive capability acquisition, Chain-of-Thought data augmentation to foster intermediate reflection, and Reinforcement Learning with Verifiable Rewards (RLVR) to further refine self-correction through reward-driven optimization. This approach achieves substantial improvements over the official challenge baselines. On the evaluation set, our best system attains a WER/CER of 11.57% for Track 1 and a tcpWER/tcpCER of 17.67% for Track 2. Comprehensive ablation studies demonstrate the effectiveness of each component under challenge constraints.
Index Terms: multilingual speech recognition, speech language model, speech diarization, self-correction, reinforcement learning.
1 Introduction
The MLC-SLM challenge focuses on multilingual conversational speech recognition and speaker diarization tasks, encompassing 11 languages: English (en), French (fr), German (de), Italian (it), Portuguese (pt), Spanish (es), Japanese (jp), Korean (ko), Russian (ru), Thai (th), and Vietnamese (vi). The English subset contains approximately 500 hours of recordings from diverse regions, including British, American, Australian, Indian, and Philippine English. Each of the remaining languages contributes around 100 hours, resulting in approximately 1,500 hours of multilingual conversational speech data.
The challenge called for participants to develop end-to-end speech language models for both ASR and speaker diarization tasks, presenting significant technical challenges in multilingual processing, conversational speech understanding, and model optimization.
This paper presents the Seewo team’s approach to the MLC-SLM challenge, with a particular focus on Track 1’s multilingual ASR task. Our proposed system achieves substantial performance improvements over the baseline [1] through systematic architectural and training innovations. The key technical contributions include:
  • A multi-stage fine-tuning pipeline that enhances the reasoning capabilities of speech language models through fine-tuning with curriculum learning strategy
  • An investigation of various reward functions for RL-based optimization of self-correction capabilities
  • An empirical study of contextual augmentation and decoding hyperparameters on ASR accuracy
2 System overview
This section details the core configuration of our system, including model architecture, data processing, computational resources and toolkits. All experiments were conducted with the same configuration.
2.1 Foundation models
Figure 1: Our model architecture reference from SLAM-ASR [2]
Our speech language model architecture follows the SLAM-ASR framework design [2]. SLAM-ASR provides a comprehensive evaluation of different encoder modules, including the Whisper family of models and other self-supervised encoders such as WavLM Large [3] and HuBERT X-Large [4].
While these self-supervised models demonstrated superior performance in monolingual settings, they were limited by English-only pretraining, rendering them unsuitable for the multilingual task. Therefore, we adopted the encoder of Whisper large-v3-Turbo [5] as our encoder for its robust multilingual capabilities and parameter efficiency.
For the decoder component, we selected Babel-9B-Chat [6] based on two key considerations. First, chat models typically outperform base pretrained models in instruction-following tasks [2]. Second, Babel-9B-Chat’s training data comprehensively covers all 11 languages required in the MLC-SLM challenge. As demonstrated in Table 7 of[6], it outperforms other 10B-size models on most multilingual benchmarks.
To bridge the encoder and decoder, we employed a learnable projection module that maps encoder outputs to the decoder embedding space. The projector module employed a hierarchical architecture consisting of a convolutional layer followed by two linear layers, designed to downsample and align the encoder’s output features with the decoder’s embedding space. This 17.32M-parameter module effectively downsamples the speech features to 10Hz (one token per 100ms), aligning the temporal resolution with the LLM’s input requirements. The whole model architecture is shown in Figure 1.
2.2 Training loss
Our training pipeline employs a two-phase approach: Supervised Fine-Tuning (SFT) followed by Reinforcement Learning with Verifiable Rewards (RLVR). This design enables systematic enhancement of the model’s ASR capabilities while maintaining stability.
In the SFT phase, we optimize the model using two objectives. First, we apply the standard causal language modeling loss to establish basic next-token prediction capabilities. Additionally, in the Chain-of-Thought (CoT) stage, we employ a weighted loss variant that assigns different weights to various completion sections [7], allowing the model to focus on critical transcription segments. Through this approach, the model learns to transcribe the speech content after reasoning.
The RLVR phase adopts the Dr. GRPO [8] method, an optimized variant of the original GRPO framework [9]. This approach improves the model performance by carefully balancing exploration of potential improvements with the maintenance of model stability through KL divergence constraints with the reference model. The verifiable rewards ensure that the model’s reasoning capabilities are enhanced in a controlled manner.
2.3 Data and augmentation
The ASR model is trained on the MLC-SLM challenge training set, as described in Section 1.
The speaker embedding model of speaker diarization pipeline is trained on the MLC-SLM challenge training set and additional open-source datasets including CN-Celeb1, CN-Celeb2 [10], VoxBlink [11] and VoxBlink2 [12], to build a robust multilingual speaker embedding model.
For evaluation purposes, we utilized the development set (approximately 4 hours per language) as our primary test set, since the challenge did not provide an official evaluation set and limited the number of submissions. All ablation data reported in this paper are based on the development set.
We applied simple additive noise and reverberation augmentation techniques [13] to enhance model robustness. For each training sample, data augmentation was applied with a probability of 0.7; when augmentation was performed, either additive noise or reverberation was randomly selected. The RIR and noise data were randomly sampled from the SLR28 dataset [14] and the MUSAN dataset [15], respectively.
2.4 Training toolkits and resources
Our training pipeline leverages several toolkits including west [16], transformers [17], accelerate [18], peft [19] , and trl [20]. All experiments were conducted on a cluster of 24 NVIDIA A800 GPUs, each equipped with 80GB of memory.
3 Training pipeline of ASR system
We employ a multi-stage training approach, where each stage addresses specific challenges encountered during model development. The latter stage is continued fine-tuning on the checkpoint of the previous one.
3.1 Stage 1: Projector training
The chat template is also following the SLAM-ASR [2]: ”USER: <S> <P> ASSISTANT: <T>”, where <S> represents speech embedding, <P> represents the prompt, which is ”Transcribe the speech above” in English for all data, and <T> represents the corresponding transcribed text.
We train the projector module (freezing all parameters except the projector module) in the first 2000 steps, checkpoint-2000 (ckpt-2000). At the end of this stage, the loss converges to around 0.35. It ensures the speech features are basically aligned with the decoder’s embedding space, and the LLM roughly follows the transcription instruction.
3.2 Stage 2: Special tokens adaptation
Following Stage 1, we observed that the model occasionally deviated from the transcription task, generating conversational responses to the speech content rather than producing transcriptions. To address this instruction-following issue, we extended several special tokens to enhance the model’s structured output capabilities, which is inspired by Qwen2.5 [21] and LaVIT [22].
We extended Babel-9B-Chat’s vocabulary with four categories of functional tokens:
  • <LANG_XX>, Language-specific tokens for precise language identification across all 11 target languages
  • <speech> </speech>, Speech segment delimiters to clearly demarcate speech input boundaries
  • <transcribe> </transcribe>, Task control tokens to enforce transcription-focused output generation
  • <think> </think>, Reasoning framework tokens to facilitate explicit intermediate reasoning
To accommodate these new tokens, we selectively unfroze the embedding layer and language model head, enabling the model to fine-tune the weights for token representations while maintaining the integrity of the pre-trained knowledge.
The enhanced prompt template follows the basic format ”USER: <S> <P> ASSISTANT: <T>”, with two key modifications from Stage 1:
  • <S>, Speech input is encapsulated within <speech> and </speech> tags to establish distinct input boundaries
  • <T>, Assistant outputs a explicit structured format as ”<LANG_XX> <transcribe> [transcription] </transcribe>
This tokenized approach significantly improved the model’s task adherence and output consistency, leading to a loss of approximately 0.2 after 4000 steps (ckpt-4000).
3.3 Stage 3: LoRA training with multilingual prompts and context
While Stage 2 improved instruction following capability, two challenges remained: (1) code-mixing between English and the target languages in transcriptions, likely due to English prompt interference, and (2) a training loss plateau at 0.2, indicating potential for further optimization.
To address these issues, we adopted LoRA [23] adaptation on both the LLM decoder and Whisper encoder. We configured the LoRA modules with rank 16 and lora_alpha 32, providing a balance between model capacity and training efficiency. Additionally, we introduced language-specific prompts in the <P> section, which were ”Transcribe the speech above” translated into each respective language.
This adaptation strategy proved effective in tackling the code-mixing issue, reducing the training loss to 0.15 after 7000 steps (ckpt-7000).
3.4 Stage 4: SFT for reflection
Despite improvements in previous stages, the model continued to exhibit typical ASR errors, such as substitutions, insertions, and deletions. Recent studies [24] [25] have shown that large language models can be trained to self-verify and self-correct. To achieve this, we first trained the model with Chain-of-Thought (CoT) data for explicitly reflecting mistakes before outputting the transcription.
The CoT prompt template maintains the basic structure ”USER: <S> <P> ASSISTANT: <T>” but introduces a reasoning part in <T> section, which follows the format: ”<LANG_XX> <think> The speech sounds like: $hypothesis_{1}$, but it might have some [error_details], let me correct it. </think> <transcribe> $hypothesis_{2}$ </transcribe>
The $hypothesis_{1}$ is generated with the Stage 3 checkpoint (ckpt-7000) by running inference on the training data. For each hypothesis, the [error_details] is generated by computing the WER/CER against the ground truth transcriptions, and specific error details are identified whenever the WER is non-zero, ensuring that at least one relevant error is included when present in the hypothesis. The ground-truth transcription is used as $hypothesis_{2}$.
We continue fine-tuning the model from ckpt-7000 of Stage 3. The first attempt resulted in a training loss of 0.12 after 9000 steps (ckpt-9000). However, despite the lower training loss, evaluation on the development set revealed a WER/CER of 35%, which was significantly worse than that of ckpt-7000. Further analysis showed that the model often produced correct transcriptions in the $hypothesis_{1}$ section while generating incorrect final outputs in the <transcribe> tags. This indicates that the model failed to effectively transfer its intermediate reasoning to the final transcription output.
To understand this counterintuitive result, we analyzed the loss function in detail. We identified a fundamental issue with the standard causal modeling loss (Eq. 1):
$L_{c}=-\frac{1}{|I_{c}|}\sum_{i\in I_{c}}\log\mathbb{P}(t_{i}|t_{<i};\theta)$ (1)
We hypothesize that the standard loss function is suboptimal for this task, primarily due to two factors: (1) the completion length $|I_{c}|$ inversely affects the loss magnitude, and (2) the <think> section, which in most cases is twice as long as the <transcribe> section, disproportionately influenced the loss. Recent studies [26, 7] have shown that the completion-to-prompt length ratio ($R_{g}$) significantly impacts model performance. In our experiments, this imbalance resulted in close loss values for stages 3 and 4, despite substantial differences in WER.
To address this, we implemented a modified version of the prompt loss token weights (PLW) strategy [7]. Unlike the original approach, we applied $PLW<1$ to the <think> section while maintaining full masking for the prompt. This modification focused the model’s attention on the <transcribe> section through the following loss calculation:
$L=\frac{\displaystyle-\sum_{i=1}^{N}w_{i}\cdot\log p_{i}}{\displaystyle\sum_{i=1}^{N}w_{i}}$ (2)
where $w_{i}=1$ for the <transcribe> section and $w_{i}=PLW$ for the <think> section.
We resumed training from ckpt-9000 with a decaying PLW schedule: starting at 1.0 and reducing to 0.1 over the first 300 steps, then maintaining 0.1. This approach achieved a final loss of 0.1 after 10000 steps (ckpt-10000).
The adoption of the PLW strategy successfully refocused the model’s learning on the transcription segment, resulting in a substantial reduction in WER/CER, as summarized in Table 1.
Table 1: WER/CER and loss of the SFT stages
Stage training loss WER/CER
ckpt -2000 0.35 24.32%
ckpt -4000 0.2 18.99%
ckpt -7000 0.15 15.48%
ckpt -9000 0.12 35.03%
ckpt -10000 0.1 13.42%
3.5 Stage 5: RLVR for reflection
While Stage 4 successfully established the CoT completion pattern, two critical issues remained: (1) the WER/CER performance still fell short of our best experimental results, and (2) the reasoning content in the <think> section lacked meaningful analysis. After looking into the model’s outputs, we identified several systematic failure modes in the reasoning process:
  • Inaccurate error type identification in the initial hypothesis analysis
  • Insufficient or missing error details in the reflection phase
  • Inability to apply correct error details to final transcription, even when accurately identified
  • Generation of non-existent errors (hallucinations) in the reasoning process
To address these issues, we implemented a series of reward functions to guide the model to generate more meaningful reasoning content with Dr.GRPO [8], which fixed length bias of original GRPO [9]. In our experiments, we rolled out 4 samples for each group, with multinomial sampling and beam search decoding. The sampling temperature is set to 0.5 to control the diversity of the samples.
Here are the reward functions we implemented:
  • $RF_{1}$: verifies the structural correctness of the output with respect to <think> and <transcribe> tags.
  • $RF_{2}$: evaluates the accuracy of $hypothesis_{2}$ with respect to the ground truth (scaled between 0 and 1).
  • $RF_{3}$: verifies whether the error types identified in the reflection match the actual errors present in $hypothesis_{1}$.
  • $RF_{4}$: verifies whether the specific error details described in the reflection correspond to the actual errors in $hypothesis_{1}$.
  • $RF_{5}$: evaluates whether $hypothesis_{2}$ achieves a lower WER/CER than $hypothesis_{1}$ with respect to the ground truth, rewarding only if $hypothesis_{2}$ is more accurate.
$RF_{1}$ and $RF_{2}$ are used to constrain the basic performance of the model, including the format of output and the accuracy of the transcription. $RF_{3}$ and $RF_{4}$ are used to guide the model to generate more accurate reflection content. $RF_{5}$ is used to guide the model to generate more accurate transcription.
We conducted experiments with different combinations of these reward functions, all post-training experiments were based on ckpt-10000. Each of the reward score was scaled to 0 to 1.0, and the final advantage was calculated by the sum of the reward scores, with a weight of 0.8 for the $RF_{2}$ and 0.5 for the others. In table 2, we show the WER/CER results of different combinations of reward functions, compared with the best model in SFT stages.
Table 2: Results of different combinations of reward functions
Reward Functions WER/CER
ckpt-10000 (without RL) 13.42%
$RF_{1}$+$RF_{2}$ 13.15%
$RF_{1}$+$RF_{2}$+$RF_{3}$ 15.44%
$RF_{1}$+$RF_{2}$+$RF_{3}$+$RF_{4}$ 16.12%
$RF_{1}$+$RF_{2}$+$RF_{5}$ 12.73%
$RF_{1}$+$RF_{5}$ 28.29%
$RF_{2}$+$RF_{5}$ 20.75%
The combination of $RF_{1}$, $RF_{2}$, and $RF_{5}$ yielded the best performance among all reward function configurations in our experiments. This approach was inspired by SCoRe [25], which employs a two-step correction mechanism to enhance final model performance. In our experiments, the reward associated with $RF_{5}$ consistently converged above 0.8, indicating that the model developed strong self-correction capabilities. However, we also observed a phenomenon of reward hacking, where $hypothesis_{1}$ progressively collapsed and became increasingly shorter after several hundred training steps. To mitigate this issue, we introduced an additional constraint requiring $hypothesis_{1}$ to be sufficiently similar to $hypothesis_{2}$ in order to receive rewards.
For the $RF_{3}$, we observed that the model’s output was getting more accurate with the error type identification. The reward of this function converged above 0.6; however, while this did not translate into improved WER/CER. The process reward did not lead to WER/CER improvement in our experiments. For the $RF_{4}$, we found that the model was not able to generate the correct error details, the reward of this function converged to 0.5. We conjectured that the reward signal was too sparse to guide the model to generate the correct error details. Until the end of the challenge, we did not find a good way to improve the self-correction ability of the model, which would be a future work.
Besides, we found that it was important to obtain the capability step by step with a curriculum learning strategy. In the attempt of $RF_{2}$+$RF_{5}$, the <think> structure format was collapsed quickly, and the rewards of $RF_{2}$ and $RF_{5}$ kept in a low level. The GRPO reinforcement learning process was trying to getting the better sampling signal, the premise is that better sampling results exists in a group. [9]
At the end of our post-training experiments, we obtained a WER/CER of 12.73% with $RF_{1}$+$RF_{2}$+$RF_{5}$, with a checkpoint called ckpt-12500.
4 Additional Experiments
During the evaluation phase of the challenge, we also found some other approaches to improve the performance of the ASR task.
4.1 Effect of Conversational Context on ASR
Inspired by GEC-RAG [27], which demonstrated that contextual information can enhance error correction in post-processing, we explored the impact of conversational context on ASR model performance in the MLC-SLM challenge. Given the conversational and long-form nature of the audio, we hypothesized that incorporating preceding utterances could improve the model’s self-correction ability. During training, we augmented each sample by appending the transcriptions of the previous two utterances in the same conversation to the user prompt. With the new prompt, we fine-tuned the model from ckpt-7000 of Stage 3 to ckpt-9000. At inference time, the model was also provided with the transcriptions of the previous two utterances as context. The results are summarized in Table 3.
Table 3: Results of training and decoding with context
Stage training loss WER/CER
ckpt -7000 0.15 15.48%
ckpt -9000 (with context) 0.14 14.30%
It should be noted that, in this experiment, the contextual transcriptions used during evaluation were taken from the ground truth of the development set, rather than generated by the model itself. As a result, the reported results represent an upper bound on the potential benefit of context, rather than the true end-to-end performance. Due to the inefficiency of our inference pipeline, which is slow for sequential utterance processing by the conversational order, we were unable to conduct large-scale, fully automatic experiments, fully automatic experiments at scale. Improving the inference efficiency and conducting stricter end-to-end evaluations will be the focus of our future work.
4.2 Different decoding methods and hyperparameters
We tried different decoding methods and hyperparameters with the best performance model, including different beam size, different decoding length, different decoding method, etc. All experiments were conducted with the best checkpoint ckpt-12500. The results are shown in table 4. The best result in our experiments is 12.73% WER/CER with Beam+Sampling method, beam size 8, new max length 180.
Table 4: Results of different decoding methods
method beam size new max length WER/CER
multinomial sampling - 150 16.72%
beam search 10 150 14.04%
Beam+Sampling 10 150 13.90%
Beam+Sampling 12 150 14.26%
Beam+Sampling 15 150 16.39%
Beam+Sampling 8 150 13.55%
Beam+Sampling 5 150 13.68%
Beam+Sampling 2 150 13.61%
Beam+Sampling 8 180 12.73%
Contrary to our initial expectations, increasing the beam size did not yield consistent performance improvements, revealing a non-monotonic relationship between beam size and model accuracy.
5 Speaker diarization pipeline
Figure 2: Pipeline of speaker diarization
Our best speaker diarization system is based on the 3D-Speaker-Toolkit [28], following the standard speaker diarization recipe without overlap detection. The system architecture comprises three key components: (1) a deep FSMN-based [29] voice activity detection (VAD) module for precise speech segment identification, (2) a ResNet-101 speaker embedding model from Wespeaker [30] for discriminative speaker representation, and (3) a spectral clustering module [31] for speaker grouping. The complete pipeline is illustrated in Figure 2.
To optimize system performance, we conducted comprehensive ablation studies evaluating multiple speaker embedding models. Our experiments included the Wespeaker ResNet-101 [30], along with pre-trained models Eres2net [32] and CAM++ [33] from ModelScope [34]. The ResNet-101 model was specifically trained on a diverse corpus comprising CnCeleb1, CnCeleb2 [10], VoxBlink [11], and VoxBlink2 [12] datasets.
A critical optimization involved adjusting the VAD module’s max_single_segment_time parameter. The default 60-second setting in 3D-Speaker-Toolkit exceeds the Whisper encoder’s maximum length capacity, which degrades the performance of the ASR task. Furthermore, our training data analysis revealed that utterances typically fall under 10 seconds. Based on this observation, we systematically evaluated different combinations of speaker embedding models and segment duration settings. Table 5 presents the comparative performance metrics, demonstrating that the ResNet-101 model with an 8-second segment duration achieves the best overall performance, yielding a 16.78% DER and 18.03% tcpWER.
Table 5: Analysis of multiple speaker embedding models with different segment time configurations
Embedding Model Max Seg.Time(s) DER tcpWER
CAM++ 60 16.71% 24.64%
CAM++ 8 17.28% 19.32%
Eres2net 8 17.22% 19.27%
ResNet101 8 16.78% 18.03%
6 Conclusions
In this work, we present a comprehensive multi-stage training pipeline for multilingual conversational ASR and speaker diarization, integrating curriculum learning, Chain-of-Thought data augmentation, and RLVR-based self-correction. Our approach achieves substantial improvements over the official challenge baselines. We summarize the results of our main version submitted to the evaluation set in table 6 and 7, to demonstrate the effectiveness of each component. Despite these advances, challenges remain in fully automating context-aware inference and further improving the granularity of self-correction. Future work will focus on enhancing inference efficiency, developing more sophisticated reward functions for reasoning, and extending our methods to broader real-world multilingual and conversational scenarios.
Table 6: Results submitted to the evaluation set of track 1
WER/CER
checkpoints method dev set test set
baseline [1] - 21.49% 20.17%
ckpt-4000 Special token SFT 18.99% 20.29%
ckpt-7000 LoRA SFT 15.48% 13.86%
ckpt-10000 CoT SFT 13.42% 12.26%
ckpt-12500 RLVF($RF_{1}$+$RF_{2}$) 13.15% 11.79%
ckpt-12500 RLVF($RF_{1}$+$RF_{2}$+$RF_{5}$) 12.73% 11.57%
Table 7: Results submitted to the evaluation set of track 2
tcpWER/tcpCER
checkpoints Max Seg.Time(s) dev set test set
baseline [1] - 76.12% 60.39%
CAM++ 60 24.64% 22.39%
CAM++ 8 19.32% 18.34%
ResNet101 8 18.03% 17.67%